Pure cyclohexane, C6H12, has a freezing point of 6.53 celcius. Its freezing point depression constant is: Kf=20.0 degrees celcius m^-1. A solution was made by taking 10.55g of an unknown nonelectrolyte and dissolving it in 125.0g of cyclohexane. The measured freezing point of the solution was -3.28 degrees Celsius. Calculate the molecular weight of the unknown substance.
Answer #1

Δ Tf = 6.53 - (-3.28) = 9.81 oC

use:

Δ Tf = Kf*mb

9.81 = 20.0 *mb

mb= 0.4905 molal

m(solvent)= 125.0 g

= 0.125 kg

use:

number of mol,

n = Molality * mass of solvent in Kg

= (0.4905 mol/Kg)*(0.125 Kg)

= 6.131*10^-2 mol

mass(solute)= 10.55 g

use:

number of mol = mass / molar mass

6.131*10^-2 mol = (10.55 g)/molar mass

molar mass = 172.1 g/mol

Answer: 172 g/mol